OpenStudy (anonymous):

hey guys , a basic question how to prove that (a^2)+(b^2) > a*b always

OpenStudy (anonymous):

(a-b)^2>=0 this implies a^2+b^2-2ab>=0 this implies a^2+b^2>=2ab>ab

OpenStudy (anonymous):

@majdnation

OpenStudy (anonymous):

can i use that to prove (a^2)+(ab)+(b^2)>=0

OpenStudy (phi):

(a^2)+(b^2) > a*b always is not true unless there are restrictions on a and b. e.g. a=b=0 gives 0>0 which is false

OpenStudy (anonymous):

If a and b are not 0 then (a-b) is not zer and when squared is always greater than 0 therefore (a-b)^2>0 that is a^2+b^2-2ab>0 that is a^2+b^2>2ab>ab

OpenStudy (anonymous):

so if b!=0 , only if a=b this will be wrong thank you all